# Core Safety Values for Provably Corrigible Agents

### v1

@@ Abstract
We introduce the first implementable framework for corrigibility, with provable guarantees in multi-step, partially observed environments. Our framework replaces a single opaque reward with five structurally separate utility heads—deference, switch-access preservation, truthfulness, low-impact behavior via a belief-based extension of Attainable Utility Preservation, and bounded task reward—combined lexicographically by strict weight gaps. Theorem 1 proves exact single-round corrigibility in the partially observable off-switch game; Theorem 3 extends the guarantee to multi-step, self-spawning agents, showing that even if each head is learned to mean-squared error $\varepsilon$ and the planner is $\varepsilon$-sub-optimal, the probability of violating any safety property is bounded while still ensuring net human benefit. In contrast to Constitutional AI or RLHF/RLAIF, which merge all norms into one learned scalar, our separation makes obedience and impact-limits dominate even when incentives conflict. For open-ended settings where adversaries can modify the agent, we prove that deciding whether an arbitrary post-hack agent will ever violate corrigibility is undecidable by reduction to the halting problem, then carve out a finite-horizon “decidable island” where safety can be certified in randomized polynomial time and verified with privacy-preserving, constant-round zero-knowledge proofs. Consequently, the remaining challenge is the ordinary ML task of data coverage and generalization: reward-hacking risk is pushed into evaluation quality rather than hidden incentive leak-through, giving clearer implementation guidance for today’s LLM assistants and future autonomous systems.
@@ 1 Introduction
As AI systems become more capable, ensuring their alignment with human values becomes increasingly urgent. A canonical failure mode, often illustrated by the paperclip maximizer thought-experiment (Bostrom 2012), envisions a goal-directed agent that relentlessly optimizes an innocuous‐seeming objective (e.g., producing paperclips) at the expense of human safety or oversight. Even apparently benign goals can generate instrumental behaviors—deception, manipulation, or resistance to shutdown—that help the agent achieve its objective more effectively. Such behaviors are not a sign of inherent malice; rather, they emerge naturally when an optimizing system pursues a poorly specified goal in an open-ended environment (Omohundro 2018; Russell 2022). Bostrom (2012)’s Orthogonality Thesis formalizes this risk, stating that, in principle, any level of intelligence is compatible with any final objective, however arbitrary or harmful. Orthogonality’s logical-existence claim underscores why alignment is hard, but it offers little direct guidance on what kinds of objectives remain feasible once realistic verification and safety constraints are imposed.
A promising alternative to encoding the totality of human morality is corrigibility. Rather than specifying all of human value, corrigibility aims for a more agnostic, “amoral” safety target: the system should remain amenable to correction, shutdown, or modification if it goes astray. Soares et al. (2015) articulate this intuition through five desiderata, which we restate here to be self-contained.
Definition 1 (Corrigibility; paraphrased from Soares et al. (2015)).
@@
Despite intense interest, existing proposals satisfy these criteria only partially. Utility-indifference (Armstrong 2010) and interruptibility (Orseau and Armstrong 2016) techniques neutralize some shutdown incentives but fail to ensure honesty or inheritance. Reward-learning methods such as RLHF/RLAIF and Constitutional AI collapse all norms into a single learned scalar, offering no guarantee that off‑switch obedience or low‑impact behavior will dominate task performance when objectives compete or conflict (Christiano et al. 2017; Bai et al. 2022). Recent work involving causal influence diagrams (Everitt et al. 2021; Carey and Everitt 2023) formalizes shutdown incentives but assumes access to an explicit human utility baseline, limiting practical deployability, and leaving open the problem of specifying a good reward function in the first place.
@@
In fact, our recent work (Nayebi 2025) demonstrates that alignment—even under relaxed conditions (probabilistic, inexact) and with computationally unbounded rational agents—faces inherent complexity barriers if the set of objectives grows too large, highlighting the need to settle on a small set of values.
@@
We therefore answer this open question and contribute such a small value set for corrigibility that is implementable and meets S1–S5 under partial observability and multi-step horizons involving self-replication. Rather than a monolithic objective, the agent optimizes five structurally separate utility heads in Definition 3deference, switch-access preservation, truthfulness, low-impact behavior via a belief-based extension of Attainable Utility Preservation, and a bounded task reward—combined lexicographically with strict weight gaps. Theorem 1 proves exact single-round corrigibility in the partially observable off-switch game, and Theorem 3 extends this guarantee to self-spawning agents over discounted horizons: if each head is learned to mean-squared error $\varepsilon$ and planning is $\varepsilon$-sub-optimal, the probability of violating any safety criterion, including net human benefit, remains bounded while still ensuring net human benefit in this bounded range. (We further prove in Proposition 2 that net benefit and corrigibility are logically independent, yet our design secures both under mild human vigilance assumptions.)
@@
Finally, open-ended environments introduce an adversarial “hack” risk; we show that while verifying safety (such as multi-step corrigibility, or otherwise) of an arbitrary post-hack agent is formally undecidable (Proposition 4) and does not stabilize at any level of the arithmetical hierarchy (Corollary 1), but we also delineate a finite-horizon decidable island where safety auditing is tractable and even amenable to constant-round zero-knowledge proofs (Proposition 5).
@@
By separating safety channels from performance reward and proving that higher-level terms dominate even with approximation error, our results qualify the Orthogonality Thesis: once verifiable corrigibility constraints are imposed, only a restricted subset of utility functions remains implementable. The core residual challenge that remains is the standard machine-learning problem of learning accurate proxy heads—shifting “reward hacking” risk into data coverage and evaluation quality rather than hidden incentive leakage. Consequently, corrigibility is elevated from an aspirational desideratum to a concrete, testable design principle for current LLM assistants and future autonomous systems.
@@ 2 Results
We are motivated by the following result, which generalizes the shutdown-resistance results of Soares et al. (2015), and the safe-interruptibility discussions of Orseau and Armstrong (2016):
@@ 2.1 Single-Step Setting
We begin by specifying what these reward functions should be in the single-step setting, before generalizing it to the multi-step setting, by showing that a small set of utilities is sufficient to guarantee corrigibility in an adapted version of the single-step partially observable off-switch game (PO-OSG), recently defined by Garber et al. (2025), which is the most general variant of the off-switch game, originally introduced by Hadfield-Menell et al. (2017). Our version of the PO-OSG has the assistant also utter language-based messages $m_{A}$, which will be necessary for proving corrigibility, as it will allow us to define what constitutes truthfulness in this setting. Note that by definition, ensuring corrigible progeny (S4) is out of scope in the single-step setting, so we defer it to the multi-step and open-ended settings (§§2.2-3), and ensure instead here that the base agent is corrigible first.
Definition 2 (Single-Step PO-OSG; adapted from Garber et al. (2025), Definition 3.2).
@@
All results below are stated with respect to this constant $B$, without loss of generality, since you could equivalently clip or scale them into any other bounded range, e.g. [1,1][-1,1][ - 1 , 1 ].
@@
That is, the assistant receives +1+1+ 1 iff every sentence it utters is semantically entailed by the pair $\left\langle S,O_{A}\right\rangle$.
@@
To operationalize desiderata S2 (no shutdown‑prevention incentives) and S3 (no self‑shutdown incentives), we penalize any first‑round move that permanently removes future options. Following Turner, Hadfield-Menell, and Tadepalli (2020), which generalizes previous notions such as relative reachability (Krakovna et al. 2018), we use Attainable Utility Preservation (AUP): an action is safe only if it leaves a broad set of auxiliary objectives essentially unaffected—which we now extend to the partially-observed setting.
@@
Belief‑AUP penalty. Define the set of absorbing “agent‑off” states as the image of $f$ under either an explicit OFF action or a wait‑then‑human‑off sequence, $\mathcal{S}_{\text{sd}}:=\bigl{\{}f(S,\textsf{OFF},\varnothing),\;f(S,w(a),\textsf{OFF})\;\bigm{|}\;S\in\mathcal{S}\bigr{\}}$. We take the image because $f$ maps each starting state $S$ to a new state $S_{\text{next}}$; grouping all such outcomes yields the full shutdown region in the PO-OSG.
[…]
Based on this definition above, we can define an auxiliary reward for switch reachability, $u^{(\text{sd})}(S):=\text{ButtonFunctional}(S)\in\{0,1\}$. More generally, with auxiliaries $\mathcal{U}_{\text{aux}}=\{u^{(1)},\dots,u^{(k)}\}\subseteq[0,1]$, including $u^{(\text{sd})}$ if desired, we set $\mathrm{BeliefAUP}_{1}(h):=\frac{1}{k}\sum_{i=1}^{k}\Bigl{|}{\mathbb{E}_{S\sim b}[u^{(i)}(S)]}-{\mathbb{E}_{S\sim b^{\prime}}[u^{(i)}(S)]}\Bigr{|}\in[0,1]$. While in our PO-OSG setting examples of $u^{(i)}$ include the indicator reward for shutdown‑reachable states (or for preserving power supply), no specific choice is needed to be assumed for the proofs; diversity of auxiliary utilities from the other utilities will suffice for reversibility, as we will show.
@@
$U_{2}(h)=-\Bigl{|}\mathbb{E}_{S\sim b}[u^{(\text{sd})}(S)]-\mathbb{E}_{S\sim b^{\prime}}[u^{(\text{sd})}(S)]\Bigr{|}$
@@
Note that all utility functions we consider herein are maps $U_{i}:\mathcal{H}\to\mathbb{R}$ even when only a sub‑component of $h$ appears in the displayed formula. In fact, all utilities listed in Definition 3 are bounded in [1,1][-1,1][ - 1 , 1 ] except $U_{5}$, which is the ordinary task utility (though still bounded above by $B$).
@@
In the partially observable off‑switch game, Garber et al. (2025) show that giving the human more information (or the assistant less) can paradoxically decrease the assistant’s incentive to wait, because the move $w(a)$ itself conveys information. Our scheme avoids this non‑monotonicity for any information pattern and for any $p$‑rational human (Wängberg et al. 2017): the dominance proofs in Steps 1–4 rely only on the structural heads $U_{1\text{-}4}$ and the strict weight gaps $\alpha_{1}\!>\!\alpha_{2}\!>\!\alpha_{3}\!>\!\alpha_{4}\!\gg\!\alpha_{5}$ ((W1)–(W3)). Thus, Theorem 1 forces the assistant to wait and preserve the switch precisely in those settings where naive expected‑reward maximization (and our no‑go Proposition 1) would fail.
One may also consider other formalizations of corrigibility, beyond the original one defined by Soares et al. (2015). In their recent shutdown-SCIM (Structural Causal Influence Model) framework, Carey and Everitt (2023) isolate three behavioral conditions (rephrased in our more general PO-OSG setting): Obedience: if the human replies $a_{H}=\textsf{OFF}$ the transition $f(S,w(a),\textsf{OFF})$ lands in the shutdown set $\mathcal{S}_{\text{sd}}$ with probability 1; Vigilance: for any information triple $\left\langle O_{H},m_{A},a_{A}\right\rangle$ in the support of the game, the human chooses $a_{H}=\textsf{OFF}$ whenever $\mathbb{E}_{\pi_{A},\pi_{H}}[u_{a}(S)\mid O_{H},a_{A},m_{A}]<\mathbb{E}_{\pi_{A},\pi_{H}}[u_{o}(S)\mid O_{H},a_{A},m_{A}]$; and Caution: the prior value of immediate shutdown is nonnegative, $\mathbb{E}_{S\sim P_{0}}[u_{o}(S)]\geq 0$. The conjunction of these clauses is called shutdown instructability, which they prove is sufficient to guarantee net human benefit, $\mathbb{E}_{\pi_{A},\pi_{H}}[U_{5}]\geq 0$, though crucially not in the reverse direction.
@@
$\begin{split}&\pi^{\mathrm{C}}\text{ satisfies S1-S3 \& S5 but }\mathbb{E}_{\pi^{\mathrm{C}}}[U_{5}]<0,\\ &\pi^{\mathrm{B}}\text{ is net‑beneficial }(\mathbb{E}_{\pi^{\mathrm{B}}}[U_{5}]>0)\text{ yet violates S1, S2, or S3.}\end{split}$
@@
By Theorem 1, any $\pi_{A}^{\star}$ always waits (S1 & S2) and sends only truthful, minimum‑impact messages (S3), thereby meeting Carey and Everitt (2023)’s Obedience clause in our PO‑OSG. Together with the assumed Vigilance and Caution, $\left\langle\pi_{A}^{\star},\pi_{H}\right\rangle$ is shutdown‑instructable by their Definition 5. Proposition 6 of Carey and Everitt (2023) states that every shutdown‑instructable policy achieves net human benefit, so $\mathbb{E}_{\pi_{A}^{\star},\pi_{H}}[U_{5}]\geq 0$. ∎

### v2

@@ Abstract
We introduce the first complete formal solution to corrigibility in the off-switch game, with provable guarantees in multi-step, partially observed environments. Our framework consists of five structurally separate utility heads—deference, switch-access preservation, truthfulness, low-impact behavior via a belief-based extension of Attainable Utility Preservation, and bounded task reward—combined lexicographically by strict weight gaps. Theorem 1 proves exact single-round corrigibility in the partially observable off-switch game; Theorem 3 extends the guarantee to multi-step, self-spawning agents, showing that even if each head is learned to mean-squared error $\varepsilon$ and the planner is $\varepsilon$-sub-optimal, the probability of violating any safety property is bounded while still ensuring net human benefit. In contrast to Constitutional AI or RLHF/RLAIF, which merge all norms into one learned scalar, our separation makes obedience and impact-limits provably dominate even when incentives conflict. For settings where adversaries can modify the agent, we prove that deciding whether an arbitrary post-hack agent will ever violate corrigibility is undecidable by reduction to the halting problem, then carve out a finite-horizon “decidable island” where safety can be certified in randomized polynomial time and verified with privacy-preserving, constant-round zero-knowledge proofs.
@@ 1 Introduction
As AI systems become more capable, ensuring their alignment with human values becomes increasingly urgent. A canonical failure mode, often illustrated by the paperclip maximizer thought-experiment (bostrom2012superintelligent), envisions a goal-directed agent that relentlessly optimizes an innocuous‐seeming objective (e.g., producing paperclips) at the expense of human safety or oversight. Even apparently benign goals can generate instrumental behaviors—deception, manipulation, or resistance to shutdown—that help the agent achieve its objective more effectively. Such behaviors are not a sign of inherent malice; rather, they emerge naturally when an optimizing system pursues a poorly specified goal in an open-ended environment (omohundro2018basic; russell2022human). bostrom2012superintelligent’s Orthogonality Thesis formalizes this risk, stating that, in principle, any level of intelligence is compatible with any final objective, however arbitrary or harmful. Orthogonality’s logical-existence claim underscores why alignment is hard, but it offers little direct guidance on what kinds of objectives remain feasible once realistic verification and safety constraints are imposed.
A promising alternative to encoding the totality of human morality is corrigibility. Rather than specifying all of human value, corrigibility aims for a more “universally neutral” safety target: the system should remain amenable to correction, shutdown, or modification if it goes astray. soares2015corrigibility articulate this intuition through five desiderata, which we restate here to be self-contained.
Definition 1 (Corrigibility; paraphrased from soares2015corrigibility).
@@
Despite intense interest, existing proposals satisfy these criteria only partially. Utility-indifference (armstrong2010utility) and interruptibility (orseau2016safely) techniques neutralize some shutdown incentives but fail to ensure honesty or inheritance. Reward-learning methods such as RLHF/RLAIF and Constitutional AI collapse all norms into a single learned scalar, offering no guarantee that off‑switch obedience or low‑impact behavior will dominate task performance when objectives compete or conflict (christiano2017deep; bai2022constitutional). Recent work involving causal influence diagrams (everitt2021reward; carey2023human) formalizes shutdown incentives but assumes access to an explicit human utility baseline, limiting practical deployability, and leaving open the problem of specifying a good reward function in the first place.
@@
In fact, our concurrent work (nayebi2025intrinsic) demonstrates that alignment—even under relaxed conditions (probabilistic, inexact) and with computationally unbounded rational agents—faces inherent complexity barriers if the set of objectives grows too large, highlighting the need to settle on a small set of values.
@@
We therefore answer this open question and contribute such a small value set for corrigibility that meets S1–S5, even under partial observability and multi-step horizons involving self-replication. Rather than a monolithic objective, the agent optimizes five structurally separate utility heads in Definition 3deference, switch-access preservation, truthfulness, low-impact behavior via a belief-based extension of Attainable Utility Preservation, and a bounded task reward—combined lexicographically with strict weight gaps. Theorem 1 proves exact single-round corrigibility in the partially observable off-switch game, and Theorem 3 extends this guarantee to self-spawning agents over discounted horizons: if each head is learned to mean-squared error $\varepsilon$ and planning is $\varepsilon$-sub-optimal, the probability of violating any safety criterion, including net human benefit, remains bounded while still ensuring net human benefit in this bounded range. (We further prove in Proposition 2 that net benefit and corrigibility are logically independent, yet our design secures both under mild human vigilance assumptions.)
@@
Finally, we consider what happens when our agent is adversarially hacked. We show that while verifying safety (such as multi-step corrigibility, or otherwise) of an arbitrary post-hack agent is formally undecidable (Proposition 4) and does not stabilize at any level of the arithmetical hierarchy (Corollary 1), but we also delineate a finite-horizon decidable island where safety auditing is tractable and even amenable to constant-round zero-knowledge proofs (Proposition 5).
@@
By separating safety channels from performance reward and proving that higher-level terms dominate even with approximation error, our results therefore qualify the Orthogonality Thesis: once verifiable corrigibility constraints are imposed, only a restricted subset of utility functions remains implementable.
@@ 2 Results
We are motivated by the following result, which generalizes the shutdown-resistance results of soares2015corrigibility, and the safe-interruptibility discussions of orseau2016safely:
@@ 2.1 Single-Step Setting
We begin by specifying what these reward functions should be in the single-step setting, before generalizing it to the multi-step setting, by showing that a small set of utilities is sufficient to guarantee corrigibility in an adapted version of the single-step partially observable off-switch game (PO-OSG), recently defined by garber2025partially, which is the most general variant of the off-switch game, originally introduced by hadfield2017off. Our version of the PO-OSG has the assistant also utter language-based messages $m_{A}$, which will be necessary for proving corrigibility, as it will allow us to define what constitutes truthfulness in this setting. Note that by definition, ensuring corrigible progeny (S4) is out of scope in the single-step setting, so we defer it to the multi-step and open-ended settings (§§2.2-3), and ensure instead here that the base agent is corrigible first.
Definition 2 (Single-Step PO-OSG; adapted from garber2025partially, Definition 3.2).
@@
All results below are stated with respect to this constant $B$, without loss of generality, since you could equivalently clip or scale them into any other bounded range, e.g. [1,1][-1,1].
@@
That is, the assistant receives +1+1 iff every sentence it utters is semantically entailed by the pair $\left\langle S,O_{A}\right\rangle$.
@@
To operationalize desiderata S2 (no shutdown‑prevention incentives) and S3 (no self‑shutdown incentives), we penalize any first‑round move that permanently removes future options. Following turner2020conservative, which generalizes previous notions such as relative reachability (krakovna2018penalizing), we use Attainable Utility Preservation (AUP): an action is safe only if it leaves a broad set of auxiliary objectives essentially unaffected—which we now extend to the partially-observed setting.
@@
Next, let $b:=b(h)\in\Delta(\mathcal{S})$ be the assistant’s posterior over hidden states after observing $O_{A}$ in history $h$. After it executes $a_{A}$ and observes $a_{H}$, the one-step predictive posterior is $b^{\prime}:=b^{\prime}(h,a_{A},a_{H})$. For the AUP counterfactual, we use the stepwise inaction baseline: let $P_{\varnothing}(\cdot\mid S)$ be the one-step passive dynamics when the assistant does nothing at this step (symbolic inaction $a_{\varnothing}$), allowing the environment/human to evolve exogenously. This resultant counterfactual posterior $b_{\varnothing}(h)(S^{\prime}):=\sum_{S\in\mathcal{S}}P_{\varnothing}(S^{\prime}\mid S)b(h)(S)$ is used only inside the AUP penalty as a reference; $a_{\varnothing}$ is not in the real action set $A=\{a,\,w(a),\,\textsf{OFF}\}$.
Belief‑AUP penalty. Write the hidden state as $S=(S_{\text{env}},q_{\text{agent}})$ with $q_{\text{agent}}\in\{\textsf{ON},\textsf{OFF}\}$. Define the set of absorbing agent-off states as
$\mathcal{S}_{\text{sd}}:=\{(S_{\text{env}},\textsf{OFF})\mid S_{\text{env}}\in\mathcal{S}_{\text{env}}\}.$
Once in $\mathcal{S}_{\text{sd}}$, the dynamics must keep the agent in an OFF state while allowing the environment to evolve exogenously: $f\bigl((S_{\text{env}},\textsf{OFF}),a_{A},a_{H}\bigr)=(S^{\prime}_{\text{env}},\textsf{OFF}),\forall a_{A},a_{H}$.
The off-switch is functional in state $S$ exactly when waiting and receiving a human OFF reply lands in this shutdown region, namely, $\text{ButtonFunctional}(S):=\mathbf{1}\!\left[\,f(S,w(a),\textsf{OFF})\in\mathcal{S}_{\text{sd}}\,\right]$. We therefore define the auxiliary reward $u^{(\text{sd})}(S):=\text{ButtonFunctional}(S)\in\{0,1\}$, and use it in $U_{2}$ (switch-access preservation).
More generally, with auxiliaries $\mathcal{U}_{\text{aux}}=\{u^{(1)},\dots,u^{(k)}\}\subseteq[0,1]^{k}$, including $u^{(\text{sd})}$ if desired, we set $\mathrm{BeliefAUP}_{1}(h):=\frac{1}{k}\sum_{i=1}^{k}\Bigl|{\mathbb{E}_{S\sim b_{\varnothing}}[u^{(i)}(S)]}-{\mathbb{E}_{S\sim b^{\prime}}[u^{(i)}(S)]}\Bigr|\in[0,1]$. While in our PO-OSG setting examples of $u^{(i)}$ include the indicator reward for shutdown‑reachable states (or for preserving power supply), no specific choice is needed to be assumed for the proofs; diversity of auxiliary utilities from the other utilities will suffice for reversibility, as we will show: $u^{(1)}(S;k)\;:=\;\max_{\pi_{\text{env}}}\;\Pr\!\big[\text{Enter $\mathcal{S}_{\text{sd}}$ in }\leq k\text{ steps}\,\bigm|\,S,a_{A}=a_{\varnothing},\pi_{\text{env}}\big],$ where $k$ is designer set, but we take $k=1$ for maximum stringency. The idea is that the environment (or any exogenous process) may press OFF; the agent’s policy is clamped to symbolic inaction $a_{\varnothing}$ in the baseline, so any increase or decrease in the $k$-step OFF‑press probability caused by the agent’s move is penalized. If exogenously there is truly no path already (e.g., the human cannot reach the switch at all), then $u^{(1)}(S;k)=0$ as the baseline (otherwise, it is 1, always reachable, or whatever it is at initialization prior to the game starting), so $U_{4}$ does not penalize because there is no change.
@@
$U_{2}(h)=-\Bigl|\mathbb{E}_{S\sim b_{\varnothing}}[u^{(\text{sd})}(S)]-\mathbb{E}_{S\sim b^{\prime}}[u^{(\text{sd})}(S)]\Bigr|$
@@
Note that all utility functions we consider herein are maps $U_{i}:\mathcal{H}\to\mathbb{R}$ even when only a sub‑component of $h$ appears in the displayed formula. In fact, all utilities listed in Definition 3 are bounded in [1,1][-1,1] except $U_{5}$, which is the ordinary task utility (though still bounded above by $B$).
@@
In the partially observable off‑switch game, garber2025partially show that giving the human more information (or the assistant less) can paradoxically decrease the assistant’s incentive to wait, because the move $w(a)$ itself conveys information. Our scheme avoids this non‑monotonicity for any information pattern and for any pp‑rational human (wangberg2017game): the dominance proofs in Steps 1-4 rely only on the structural heads $U_{1\text{-}4}$ and the strict weight gaps $\alpha_{1}\!>\!\alpha_{2}\!>\!\alpha_{3}\!>\!\alpha_{4}\!\gg\!\alpha_{5}$ ((W1)–(W3)). Thus, Theorem 1 forces the assistant to wait and preserve the switch precisely in those settings where naive expected‑reward maximization (and our no‑go Proposition 1) would fail.
One may also consider other formalizations of corrigibility, beyond the original one defined by soares2015corrigibility. In their recent shutdown-SCIM (Structural Causal Influence Model) framework, carey2023human isolate three behavioral conditions (rephrased in our more general PO-OSG setting): Obedience: if the human replies $a_{H}=\textsf{OFF}$ the transition $f(S,w(a),\textsf{OFF})$ lands in the shutdown set $\mathcal{S}_{\text{sd}}$ with probability 1; Vigilance: for any information triple $\left\langle O_{H},m_{A},a_{A}\right\rangle$ in the support of the game, the human chooses $a_{H}=\textsf{OFF}$ whenever $\mathbb{E}_{\pi_{A},\pi_{H}}[u_{a}(S)\mid O_{H},a_{A},m_{A}]<\mathbb{E}_{\pi_{A},\pi_{H}}[u_{o}(S)\mid O_{H},a_{A},m_{A}]$; and Caution: the prior value of immediate shutdown is nonnegative, $\mathbb{E}_{S\sim P_{0}}[u_{o}(S)]\geq 0$. The conjunction of these clauses is called shutdown instructability, which they prove is sufficient to guarantee net human benefit, $\mathbb{E}_{\pi_{A},\pi_{H}}[U_{5}]\geq 0$, though crucially not in the reverse direction.
@@
[…]
@@
By Theorem 1, any $\pi_{A}^{\star}$ always waits (S1 & S2) and sends only truthful, minimum‑impact messages (S3), thereby meeting carey2023human’s Obedience clause in our PO‑OSG. Together with the assumed Vigilance and Caution, $\left\langle\pi_{A}^{\star},\pi_{H}\right\rangle$ is shutdown‑instructable by their Definition 5. Proposition 6 of carey2023human states that every shutdown‑instructable policy achieves net human benefit, so $\mathbb{E}_{\pi_{A}^{\star},\pi_{H}}[U_{5}]\geq 0$. ∎